### v2

# Towards Multi-Agent Reasoning Systems for Collaborative Expertise Delegation: An Exploratory Design Study

Abstract
Designing effective collaboration structure for multi-agent LLM systems to enhance collective reasoning is crucial yet remains under-explored. In this paper, we systematically investigate how collaborative reasoning performance is affected by three key design dimensions: (1) Expertise-Domain Alignment, (2) Collaboration Paradigm (structured workflow vs. diversity-driven integration), and (3) System Scale. Our findings reveal that expertise alignment benefits are highly domain-contingent, proving most effective for contextual reasoning tasks. Furthermore, collaboration focused on integrating diverse knowledge consistently outperforms rigid task decomposition. Finally, we empirically explore the impact of scaling the multi-agent system with expertise specialization and study the computational trade off, highlighting the need for more efficient communication protocol design. This work provides concrete guidelines for configuring specialized multi-agent system and identifies critical architectural trade-offs and bottlenecks for scalable multi-agent reasoning. The code will be made available upon acceptance.
Towards Multi-Agent Reasoning Systems for Collaborative Expertise Delegation: An Exploratory Design Study
Baixuan Xu1††thanks: Equal Contribution, Chunyang Li1∗, Weiqi Wang1∗, Wei Fan1, Tianshi Zheng1, Haochen Shi1, Tao Fan1,2, Yangqiu Song1, Qiang Yang3 1The Hong Kong University of Science and Technology 2WeBank 3The Hong Kong Polytechnic University {bxuan, cliei, wwangbw}@cse.ust.hk
1 Introduction
Collective intelligence, the emergent problem-solving capability arising from structured group interactions, has long been recognized as a cornerstone of complex human decision-making Surowiecki (2004). Through mechanisms like deliberative debate and systematic knowledge integration, human collectives consistently outperform individual experts in tasks requiring multi-perspective analysis and contextual synthesis.
The recent evolution of large language and reasoning models (LLMs/LRMs) Yang et al. (2025); Jaech et al. (2024); Team et al. (2025) has spurred parallel investigations into machine collective intelligence. Contemporary research has developed artificial analogs of human collaboration patterns through techniques such as multi-agent debate frameworks and workflow orchestration Li et al. (2024c); Du et al. (2024); Liang et al. (2024). These approaches primarily focus on either enhancing individual model performance through collective verification processes or establishing general-purpose problem-solving workflow pipelines.
A prevalent strategy to enhance collective intelligence in these systems is collaborative expertise specialization, where LLMs are instructed to simulate specific expert personas (e.g., “act as an experienced lawyer”) Li et al. (2024a); Xu et al. (2024a). This approach is hypothesize to operate through two primary mechanisms: (1) Knowledge Recall, activating relevant domain-specific knowledge latent within the LLM via contextual role framing, and (2) Perspective Synthesis, leveraging diverse expert viewpoints to foster emergent, robust problem-solving patterns.
Although expertise specialization is widely adopted in multi-agent system Wang et al. (2024a); Li et al. (2024a), the impact of varying collaborative expertise domain on distinct downstream scenarios remains underexplored. To address this gap, we empirically evaluate the influence of different collaborative expertise configurations on task performance across four representative domains from MMLU-pro Wang et al. (2024d). Our findings demonstrate a positive correlation between task performance and the alignment of group expertise with the task domain, providing crucial insights for optimizing expertise specialization in multi-agent systems.
Another dimension concerning the multi-agent system design is the collaboration paradigm—namely, the mechanism governing how specialist agents interact in a multi-agent system. Currently, the collaboration paradigm predominantly used in recent studies could be categorized into two kinds: (1) Diversity-Driven Perspective Integration, where agents, often embodying different viewpoints or roles, are encouraged to generate diverse responses to enrich the solution space Wang et al. (2024b); Chen et al. (2024b); Hu et al. (2025). (2) Structured Workflow Cooperation, where different agents are assigned distinct sub-tasks within a predefined pipeline to collaboratively construct a solution Chen et al. (2024c); Hong et al. (2024). To understand the preference of collaboration paradigm in collaborative expertise specialization, we design comparative experiments which unveil the performance differences between paradigms. Our observations reveal a consistent advantage for diversity-driven collaboration over structured workflow collaboration, suggesting the superiority of the diversity-driven paradigm. Detailed analysis regarding viewpoint diversity are also conducted to study the impact of agent response diversity in multi-agent system, where a higher diversity could indicate a better performance.
Finally, constructing large-scale multi-agent system has become a critical, yet often enigmatic aspect of multi-agent system design Chen et al. (2024c); Piao et al. (2025). While intuition and some preliminary studies Qian et al. (2024); Li et al. (2023) suggest that larger groups would lead to a better reasoning performance, the actual effectiveness of scaling within the context of collaborative expertise specialization and the potential computation-performance trade-off, are not well understood. Motivated by the lack of clarity on scaling effects in collaborative expertise specialization, we specifically study how performance scales in multi-agent systems composed of specialized experts. Our systematic experiments involved incrementally increasing the system scale to examine potential scaling laws. The results uncover non-linear dynamics; specifically, adding more experts tends to improve the collective reasoning ability of the system. This positive trend holds regardless of whether the larger system scale contains greater viewpoint diversity or a more comprehensive workflow structure, indicating a general benefit to increasing the number of expert agents and encouraging such designs for enhanced system performance. Furthermore, our analysis of the computational trade-offs associated with system scaling reveals that, while the system would benefit from the expansion, there remains a critical need for more efficient communication protocols between agents for more scalable and cost-effective multi-agent reasoning.
2 Related Works
2.1 Multi-Agent Collaboration
Multi-Agent Collaboration adopts multiple LLMs to solve the problem collaboratively. Abundant researches have investigated the multi-agent collaboration framework to improve decision-making capability of the system Wang et al. (2024b); Liang et al. (2024); Du et al. (2024). In addition to collaboration among LLMs, several researchers instruct the agents to cooperate in a workflow to study the multi-agent systems’ ability of solving real world challenges Li et al. (2024b); Xu et al. (2024b); Chen et al. (2024a). While Qian et al. (2024), Yang et al. (2024) and Wang et al. (2024c) has investigate the effect of varying the scale of multi-agent system on reasoning and simulation, prior researches have not systematically examined the interplay between collective expertise specialization, collaboration mechanisms, and the impact of system scale simultaneously. In this work, we conduct extensive experiments to formally analyze the influence of these three critical dimensions on multi-agent collaborative reasoning. Our findings provide actionable insights toward more effective system design.
2.2 LLMs as Domain Experts
The rapid evolution of large language models (LLMs) has endowed them with vast repositories of domain-specific knowledge, enabling their application across a wide range of expert tasks. Recent research has explored the potential of LLMs to emulate specific personas by conditioning them on detailed character profiles Chan et al. (2024); Samuel et al. (2024); Xu et al. (2023). These studies demonstrate that by providing LLMs with demographic or role-specific prompts, they can effectively exhibit human-like personality traits and behaviors. Furthermore, Kong et al. (2024) and Xu et al. (2023) have shown that instructing LLMs to simulate domain experts can enhance their reasoning capabilities in specialized contexts. Despite these advancements and the growing prominence of multi-agent systems in research, the specific impact of collaborative expertise specialization on reasoning performance remains underexplored. In this paper, through meticulously designed experiments, we systematically investigate the impact of expertise specialization within multi-agent reasoning systems. Our findings reveal that simulating specialized roles significantly enhances performance on tasks requiring contextual reasoning, while showing limited influence on those primarily dependent on factual recall or mathematical deduction.
3 Preliminary
3.1 Problem Setup
Formally, given a multi-agent system where n indicates the number of agents inside the system and represents the i-th agent of the system, a query , and a set of candidate options . A multi-agent system reasoning process could be expressed as:
where stands for the final answer generated by the system. represents the answer of agent i, stands for the communication protocol manually customized by the design of the system which aggregate the answer of each agents into the final answer. Typically, it could be majority vote, debate, etc Kaesberg et al. (2025); Liu et al. (2024a). In our specific setup, we adopt a sequential processing communication mechanism inspired by Qian et al. (2024) to prevent context explosion Liu et al. (2024b); Xu et al. (2024c). In this mechanism, for , agent receives the complete output generated by the immediately preceding agent . In contrast, from the preceding agents {}, receives only the final answers. The detailed communication algorithm could be found in Appendix A algorithm 1.
3.2 Dataset
For our experiments, we select four distinct domains from MMLU-pro Wang et al. (2024d): Math, Health, Business, and Law. These four domains were selected for being representative and frequently studied in contemporary multi-agent reasoning research Cui et al. (2023); Lei et al. (2024); Ghezloo et al. (2025). We further classify these four domains into three categories based on the primary reasoning type required: (1) Mathematical Reasoning: Domains requiring formal mathematical deduction to derive the answer. (2) Factual Recall Reasoning: Domains primarily requiring the recall of domain-specific factual knowledge, seldom needing extensive reasoning steps. (3) Contextual Reasoning: Domains requiring not only the retrieval of relevant expert knowledge but also its application within the reasoning process of specific scenarios or contexts. This choice of evaluation domains and fine-grained classification of their reasoning types allow us to investigate the effects of collaborative expertise specialization on multi-agent system from a more systematic manner.
3.3 Collaborative Expertise Specialization
In this paper, we primarily studied the effect of collaborative expertise specialization on better multi-agent system design from the perspective of expert-domain alignment, collaboration paradigms and system scale. To formalize the role and responsibility of the agents in the multi-agent system, we define each expert to be of the following format:
Where stands for the agent i, EG, FR and R represent Expert Group, Formal Role, Responsibility respectively. ID represents an agent’s index within the group of all agents who share the same specific expert role.
3.4 General Experiment Setup
As detailed in Section 3.2, we select 4 representative domains from MMLU-pro to investigate the effects of collaborative expertise specialization. To be consistent with all experiments, we utilize DeepSeek-R1-Distill-Qwen-7B DeepSeek-AI et al. (2025) as the foundational model for all experiments. Each agent is initialized with its specific expert description and responsibilities via its system prompt, while the task instance is provided through the user prompt. The detailed prompts could be found in Appendix B. All experiments adopt accuracy as the evaluation metric.
4 Leveraging with the “Right” Agent
Expertise specialization is a widely adopted technique in agent research, demonstrably enhancing the reasoning capabilities of Large Language Models (LLMs) within specific domains Li et al. (2024b). While the benefits of specialization for individual agents are well-established, the effect of collaborative expertise specialization on the collective reasoning performance of multi-agent systems remains underexplored. This section presents our experimental investigation into this critical area, designed to unveil how different collaborative specialization configurations influence the reasoning capabilities of multi-agent systems.
4.1 Setup
Considering the primary principle of multi-agent reasoning system is to incorporate more diverse agent viewpoints and integrate them in the final answer Liang et al. (2024), in our experiments, we adopt diversity-driven collaboration paradigm where we distribute each agent with a specific domain expert configuration and instruct them to generate responses based on their expertise. At this stage, we fix the size of the multi-agent reasoning system to be 3 for controllable computational cost. We employ GPT-4o OpenAI et al. (2024). The model is prompted with the target expert domain, the group size as inputs. The detailed prompts utilized for this automated role generation process are provided in Appendix C.
4.2 “Right” Expertise Helps Reasoning
Our experiments demonstrate a clear performance advantage when the collaborative expertise specialization of the multi-agent system aligns with the domains of the downstream task. Misaligned expertise configurations often underperform compared to aligned ones. This primary finding is quantitatively supported by the results presented in Table 1. Specifically, in 75% of the aligned cases (diagonal entries), the system achieved the highest accuracy compared to configurations where the agent group simulates expertise from other domains for the same task.
To gain a more nuanced understanding of when expertise alignment is most beneficial, we analyzed performance according to the primary reasoning type required by each domain, as categorized in Section 3.2. Our analysis reveals that the benefits of expertise alignment are most pronounced for tasks demanding contextual reasoning. Systems operating on Health and Law domains exhibits an average relative performance improvement of 6.75% when expertise was correctly aligned, compared to the misaligned configurations which perform the second best for those tasks. Conversely, for domains requiring mathematical reasoning (Math/Business), the specialized experts yield only marginal gains or even degradation relative to misaligned configurations. We hypothesize this divergence stems from the inherent strengths of LLMs on math. These models often possess robust mathematical reasoning capabilities due to extensive pre-training, potentially reducing the added value of specialized agents. Contextual reasoning tasks, however, appear to benefit more from the structured integration of specialized perspectives provided by the multi-agent reasoning system since applying domain knowledge in these contexts often requires nuanced interpretation, synthesis of information, and reasoning beyond direct mathematical deduction.
| Dom.Exp. | Math | Fina | Med | Law | ||
| Math | 78.0 | 76.3 | 76.3 | 76.4 | 2.1% | |
| Business | 65.4 | 64.3 | 62.4 | 62.4 | -1.7% | |
| Health | 28.9 | 26.8 | 30.4 | 26.1 | 5.2% | |
| Law | 18.3 | 19.2 | 18.5 | 20.8 | 8.3% |
4.3 Analysis on Expert-Domain Alignment
Furthermore, our experimental results reveal a positive correlation between how well the simulated group expertise aligns with (or matches) the downstream task domain and the observed performance gain. This relationship is visualized in the task-domain relevance analysis presented in Figure 3. Specifically, configurations where the simulated expertise is more relevant to the target task domain tend to yield greater performance improvements compared to less relevant configurations.
To quantify this expertise-task relevance, we first establish a relevance matrix. We randomly sample 100 instances from each of the four primary task domains. For each instance, we prompt Deepseek-V3 DeepSeek-AI et al. (2025) to identify a list of 2-3 key expertise domains pertinent to solving the task. We then aggregate these identified candidate domains across all instances within each primary task domain. The relevance scores are calculated by counting the occurrences where a specific knowledge domain (e.g., Business) is deemed relevant for tasks in a primary domain (e.g., Math). These frequencies form a relevance matrix, visualized as a heatmap in Figure 3, where deeper color indicate higher relevance scores.
Comparing this relevance heatmap with the performance results, we observe a consistent pattern supporting our initial finding—Higher expertise-domain relevance, indicated by deeper colors in the heatmap entries, generally corresponds to better reasoning performance. Many cells with high relevance scores in Figure 3 correspond to performance that are bolded or underlined in Table 1, signifying the best or second-best performance among group expertise specialization performance for that task domain. Conversely, low relevance scores typically correspond to misaligned configurations which barely demonstrate distinct advantages conferred by their specific (misaligned) expertise.
Our findings further support the established use of collective expertise specialization in multi-agent reasoning systems, while simultaneously highlighting the critical importance of aligning expertise design with the specific requirements of the target downstream domains, paving a fundamental guidance for future specialization technique application in multi-agent reasoning system design.
5 Structured Collaboration versus Diversified Discussion
A further critical consideration in multi-agent system design is the selection of an effective collaboration paradigm. Even when individual agents possess appropriate domain knowledge, the overarching mechanism governing their interaction can impact overall system performance. In this section, we present comparative experiments designed to analyze these distinct collaboration paradigms. Our objective is to investigate their potential advantages, thereby providing empirically grounded insights to guide the design of more effective multi-agent system.
5.1 Setup
Our analysis leverages the results presented in Figure 4, where we demonstrate both domain-wise and group-wise comparisons for a comprehensive overview.
The distinction between paradigms are illustrated as follows:
Diversity-Driven Collaboration:
This paradigm emphasizes assigning agents highly specialized, fine-grained expertise within a broader domain (e.g., specific sub-fields of Laws).
The objective is to foster collaboration through the integration of diverse, complementary knowledge perspectives during the reasoning process.
Each agent contributes deep expertise from a narrow viewpoint.
Structured Workflow Collaboration:
Conversely, this paradigm assigns roles based on distinct functional responsibilities within a predefined problem-solving process, in our case, solver, critic and coordinator.
Collaboration centers on agents executing specific steps and refining intermediate outputs based on their functional role, rather than primarily contributing unique domain knowledge specializations.
The differentiation between agents stems from their function within the workflow.
To ensure a plausible, accurate generation of expert role descriptions, we continue to employ GPT-4o with collaboration paradigm as extra input.
5.2 Diversity Matters in Discussion
Our primary finding is that the diversity-driven paradigm generally yields superior performance compared to the structured workflow paradigm. This advantage holds true both when considering performance from both domain-wise and group-wise perspectives.
A domain-wise analysis, depicted in Figure 4, confirms this trend. Irrespective of the domain’s primary reasoning type categorized in section 3.2, the diversity-driven approach consistently results in performance gains over structured workflow. Notably, the most substantial improvements are observed in business and health domains, which demonstrate an average relative performance increase of 1.75% under diversity-driven paradigm. This indicates the potential of expertise with finer-granularity perform well across different domains.
Examining the results from group-wise perspective further supports this conclusion. With the exception of math expert group, all other specialized groups achieve higher average performance across all task domains when employing diversity-driven paradigm. When including the math group, the overall average relative performance improvement facilitated by the diversity-driven approach across all groups was 1.25%, indicating consistent benefits regardless of the task domain encountered.
Synthesizing these observations, the diversity-driven collaboration paradigm demonstrates a consistent performance advantage over structured workflow collaboration paradigm across both different tested domains and distinct expertise configurations. This suggests that multi-agent systems could benefit significantly from collaboration structures that emphasize fine-grained expertise allocation which stimulates viewpoint diversity, providing an empirical basis for future research directions in designing multi-agent reasoning system’s collaboration pattern.
5.3 Analysis on Response Diversity
To quantitatively characterize how the collaboration paradigm influences the diversity of agent contributions, we further design a response diversity analysis. We leverage semantic embeddings derived from Sentence-BERT Reimers and Gurevych (2019). For each task instance solved by the multi-agent system, we generate embeddings for the output of each agent. We then measure the internal diversity of the system’s responses by calculating the pairwise cosine similarity between the embeddings of outputs from different agents. This provides a measure of how semantically distinct the contributions are at different stages.
The distributions of these pairwise similarity scores for both the diversity-driven and structured workflow paradigms are presented in Figure 5. The results clearly indicate that, the pairwise cosine similarity values are consistently lower for the diversity-driven collaboration paradigm compared to the structured workflow paradigm. This finding demonstrates that the diversity-driven approach, which emphasizes fine-grained expertise, fosters greater semantic diversity among agent responses throughout the collaborative reasoning, further confirming that enhancing the diversity of perspectives within a multi-agent system would be a key factor in improving its overall reasoning performance.
6 Scaling Up Reasoning Experts
Finally, the most complicated dimension in designing multi-agent systems to foster collective intelligence is the system scale. While the deployment of large-scale multi-agent systems for simulating social behaviors has received considerable attention, the implications of scaling under collaborative expertise specialization setup remain unexplored. This section details our investigation into the effects of varying system scale on both the reasoning performance of multi-agent systems and the associated computational trade-offs. We aim to elucidate how increasing the number of agents influences collective reasoning efficacy and to call for a better communication protocol design through our trade-off analysis.
6.1 Setup
We expand our experimental setup from 3 agent to systems comprising 6 and 10 agents. For these larger system sizes, we systematically replicate the experiments previously introduced, allowing for a direct comparison across different scales.
Generating coherent and appropriately specialized expert role configurations for these larger systems requires extending the initial configurations of the 3 agent system and we continue to leverage GPT-4o for this purpose. The detailed prompts employed for this role augmentation process are provided in Appendix C
6.2 More Experts, More Intelligent System
We evaluate the effect of system scale on reasoning performance by comparing the results from larger agent systems against the baseline 3 agent system. Specifically, we calculate the domain-wise relative performance difference for the system size of 6 and 10 with respect to system of size 3. These relative performance differences are illustrated in Figure 6.
Our findings reveal a consistent trend: increasing the number of agents generally enhances the multi-agent system’s reasoning performance across the evaluated domains, regardless of whether diversity-driven or structured workflow paradigm was employed. However, the magnitude of this improvement varies significantly by domain. Corroborating our earlier observations regarding domain-specific analysis in section 4, the performance gains within math domain were marginal, even when scaling up to 10 agents. Conversely, domains that necessitate substantial contextual reasoning and knowledge application demonstrate significantly larger performance improvements with increased system scale. This disparity suggests that the benefits derived from incorporating additional agents are most pronounced for tasks requiring the integration of diverse knowledge perspectives or complex, case-specific analysis inherent in non-mathematical reasoning. For domains characterized by intense mathematical reasoning, simply increasing the number of agents could barely yield diminishing returns. We believe our finding offers valuable insight for constructing large-scale multi-agent systems intended for diverse domains.
6.3 Token-Performance Trade-off
We further explore the token-performance trade-off inherent in scaling multi-agent reasoning systems by calculating the ratio of performance improvement over token overhead (PoT) with qualitative results presented in Figure 7. We use the sum of reasoning token and answer token for the calculation of token overhead and all the performance improvement and token consumption overhead are counted against system of size 3.
Our analysis reveals distinct trends both across and within domains. Cross-domain comparisons demonstrate that tasks requiring substantial contextual reasoning, such as those in health and law, yield higher PoT ratios. This suggests that increasing agent collaboration is particularly beneficial in these areas, as greater token consumption during the reasoning process leads to higher performance improvements. Conversely, mathematical reasoning tasks exhibit only marginal performance gains with additional agents, which implies smaller ensembles can achieve comparable performance with lower computational overhead, making large-scale multi-agent systems unnecessary for these tasks.
For intra-domain analysis, while structured workflows improved PoT in 75% of domains and diversity-driven approaches in 50% respectively, the critical finding is that neither collaboration paradigm guarantees a enhanced PoT across all domains tested. This widespread inconsistency in scaling behavior, regardless of the collaboration paradigm, highlights the pressing need for advancements in multi-agent communication protocols to achieve more stable and predictable performance enhancements as system complexity increases.
7 Conclusions
In conclusion, this paper systematically investigate the impact of multi-agent system expertise specialization on collective reasoning intelligence across three key dimensions: expertise-domain alignment, collaboration paradigm, and system scale. Our empirical results demonstrate that the effectiveness of aligning agent expertise with task domains is highly contingent on the reasoning type required. Furthermore, we find that collaboration paradigms emphasizing the integration of diverse perspectives outperform structured task decomposition, with response diversity playing a crucial role. Finally, a scaling law behind collaborative expertise specialization in multi-agent system is observed where stacking more experts could result in better reasoning performance. These findings provide actionable insights for designing specialized multi-agent reasoning systems in future researches and underscore the critical need for developing more efficient coordination protocol as systems scale.
Limitations
The limitations of the paper concern following several dimensions. 1. Since we focus on the reasoning capability of multi-agent system, we adopt MMLU-pro as our evaluation dataset which consist of massive reasoning tasks instead of tasks like coding. 2. To be closer to real-world scenario, we utilize only four domains (Math, Business, Health and Law) which are most focused by mainstream researches while leaving other domains underexplored. 3. To simplify the research setup and to yield more stable conclusion, we conduct our experiments on only one message propagation mechanism while overlook the potential impact could be brought by other communication mechanisms. 4. For controllable computational cost, we select DeepSeek-R1-Distilled-Qwen-7B as our base model for all the experiments, leaving multi-agent system design for larger model for future studies.
Ethics Statement
Our study involves publicly available datasets and use Large Language Models through APIs.
Consequently, the ethical considerations of this paper could be listed as follow:
Datasets: We use publicly available datasets only for academic research purpose.
We guarantee no personal data has been involved.
LLMs API: Our application of LLMs conform API provider’s policy strictly, maintaining fair use and respecting intellectual property.
Transparency: We provide detailed descriptions of our method and the prompts used in our experiments, in line with standard practices in the research community.
We will also make our code publicly available upon acceptance.
Appendix A Agent Communication Algorithm
In this section, we provide our detailed algorithm for inter-agent communication protocol and its corresponding notation table in below.
| Symbol | Meaning |
| The output without rationale of agent | |
| Full output with rationale of agent | |
| Input question | |
| The candidate answers of the question | |
| The final answer of the system |
Appendix B Role System Prompt
Appendix C Expert Generation Prompts
C.1 Primary Expert Generation Prompts
C.2 Expert Augmentation Process
Appendix D Social Group Role Examples
Appendix E Relevance Prompt
Appendix F All Experiments